# RDB2G-Bench: A Comprehensive Benchmark for Automatic Graph Modeling of Relational Databases

### v1

@@ Abstract
Relational databases (RDBs) are composed of interconnected tables, where relationships between them are defined through foreign keys. Recent research on applying machine learning to RDBs has explored graph-based representations of RDBs, where rows of tables are modeled as nodes, and foreign key relationships are modeled as edges. RDB-to-graph modeling helps capture cross-table dependencies, ultimately leading to enhanced performance across diverse tasks. However, there are numerous ways to model RDBs as graphs, and performance varies significantly depending on the chosen graph model. In our analysis, applying a common heuristic rule for graph modeling leads to up to a 10% drop in performance compared to the best-performing graph model, which remains non-trivial to identify.
To foster research on intelligent RDB-to-graph modeling, we introduce RDB2G-Bench, the first benchmark framework for evaluating such methods. We construct extensive datasets covering 5 real-world RDBs and 12 predictive tasks, resulting in around 50k graph-performance pairs for efficient and reproducible evaluations. Thanks to our precomputed datasets, we were able to benchmark 9 automatic RDB-to-graph modeling methods on the $12$ tasks over 600$\times$ faster than on-the-fly evaluation, which requires repeated model training. Our analysis of the datasets and benchmark results reveals key structural patterns affecting graph model effectiveness, along with practical implications for effective graph modeling.
@@ 1 Introduction
A relational database (RDB) is a collection of data organized into multiple tables connected by shared keys. RDBs enable systematic and efficient management of related information and have been adopted across diverse domains, including finance [17], healthcare [33], and e-commerce [29]. This widespread use has led to the emergence of various machine learning problems defined on RDBs.
@@
One important consideration in graph-based approaches is the variety of ways to model RDBs as graphs. For example, a single table row representing a user transaction on an item can be modeled either as (1) a node representing the transaction or (2) an edge linking the corresponding user and item. Additionally, one may choose to include only a subset of tables (i.e., nodes) and FK relationships (i.e., edges), excluding those considered less relevant.
@@
Despite the importance of intelligent RDB-to-graph modeling, research in this area remains in its early stages, with one main reason being the difficulty of evaluation. First, there are numerous possible graph models (see Figure 2(a) for examples from real-world RDBs), making exhaustive comparisons extremely expensive. Moreover, evaluating even a single model is computationally costly, as it typically involves training graph neural networks (GNNs) on a large graph.
@@
  • Benchmarks. We present the extensive benchmark results of 9 automatic RDB-to-graph modeling methods across 12 tasks. Thanks to our precomputed datasets, we were able to obtain the benchmark evaluation results over 600× faster than on-the-fly evaluation, which requires repeated GNN training.
@@ 2.1 Relational Deep Learning
Early approaches to relational database (RDB) learning relied on feature engineering [18] (e.g., table joins), transforming an RDB into a single table to leverage tabular machine learning [5, 20, 2, 12]. Recent works on relational deep learning (i.e., deep learning on RDBs) [8, 30, 10, 11] pioneered modeling RDBs as graphs, which facilitate the capture of cross-table dependencies. Building on these graph models, advanced methods [6, 40] have achieved improved predictive performance on RDBs.
@@ 2.2 Benchmarks on Relational Deep Learning
Distinctive property of our benchmark: Existing benchmarks provide RDBs and predictive tasks to compare machine learning (ML) methods, which are typically based on the same graph presentation of the RDBs. Our benchmark, RDB2G-Bench, however, is designed to evaluate RDB-to-graph modeling strategies, which provide graph models based on which various ML methods perform.
Specifically, by offering precomputed performance metrics for 50k graph models, RDB2G-Bench enables comparing various graph modeling strategies without repeated ML training (see Section 5.2 for efficiency gains).
@@ 2.3 Graph-based Modeling of RDBs
Our benchmark, RDB2G-Bench, enables the evaluation of various RDB-to-graph modeling methods, including both heuristic and LLM-based approaches (see Section 5.1 for the nine methods included in the benchmark), ultimately fostering the development of more advanced modeling strategies.
@@ 3 Dataset Design for RDB2G-Bench
In this section, we introduce the datasets provided in RDB2G-Bench, our RDB-to-graph modeling benchmark. Each dataset comprises graph models (i.e., representations) of a relational database (RDB) paired with downstream task performance metrics of machine learning (ML) models trained on those graph models. We first describe the considered space of graph models (Section 3.1), followed by the procedure for obtaining ML performance metrics on each graph model (Section 3.2).
@@ 3.1 Construction of Graph Models
As discussed in Section 1, there are various ways to model each RDB as a graph, and the choice of graph can significantly impact the performance of ML models built on these graphs. In RDB2G-Bench, we consider the design choices in the following two steps of RDB-to-graph modeling, both of which have a significant impact on the downstream-task performance of ML models, as shown in Section 4:
@@
That is, for each RDB (e.g., rel-avito), our dataset includes graph models that result from different combinations of design choices in Step 1 and Step 2 (refer to Figure 1 for examples). These graph models are used to train graph neural networks (GNNs) for predictive tasks (e.g., user-clicks) defined on the RDB (as described in Section 3.2), and to be used for this purpose, the graph models must satisfy several constraints. First, a valid graph model must select the task table, which the considered predictive task is defined on. Second, all selected tables should be connected to the task table via a path whose length does not exceed the number of GNN layers. Violating this constraint would result in some nodes being unreachable during message passing, leading to degraded performance. Third, (the rows of) a table can be modeled as edges only if the table has exactly two FKs and its primary key (PK) is not referenced by any FKs in other tables. Note that, for a table with more than two FKs, hyperedge modeling would be required, which is beyond our scope, and edge modeling of a table whose PK is referenced by FK renders those FK relationships unrepresentable. Also note that these constraints may lead to different graph model spaces for downstream tasks defined on the same RDB. As summarized in Figure 2(a), we constructed about 50k graph models in total, for which we collected downstream task performance metrics, as described below.
(a) Summary of the RDB2G-Bench datasets, which cover 50k graph models in total. RDB Task Name Type # Tables # Graph Performance Statistics Models Best AR2N [28] Worst rel-avito user-clicks (UC) classification 8 944 67.93 64.66 60.89 user-visits (UV) classification 944 66.33 65.97 59.83 ad-ctr (AC) regression 1304 0.039 0.040 0.044 user-ad-visit (UAV) recommendation 909 3.682 3.661 0.159 rel-event user-repeat (UR) classification 5 214 82.29 77.65 63.96 user-ignore (UI) classification 214 82.82 82.22 74.29 user-attendance (UA) regression 214 0.237 0.244 0.266 rel-f1 driver-dnf (DD) classification 9 722 74.56 73.14 67.40 driver-top3 (DT) classification 722 81.88 78.11 75.37 driver-position (DP) regression 722 3.831 3.913 4.171 rel-stack post-post-related (PPL) recommendation 7 7979 12.04 10.82 0.006 rel-trial study-outcome (SO) classification 15 36863 70.91 68.09 62.85
(b) Visualization of the RDB2G-Bench datasets on three predictive tasks. Refer to caption
@@
Figure 2: (a) We summarize the RDBs, tasks, and their associated graph models. For each classification, regression, and recommendation task, we collect AUC-ROC (%), MAE, and MAP (%), respectively, on each graph model. For each task, we report the performances on the best graph model, the worst model, and that given by AR2N modeling [28]. (b) For three tasks (driver-top3, user-attendance, post-post-related), we visualize the distribution of performances (Y-axis) across all graph models, along with training time (X-axis) and GNN parameter size (indicated by color). Note that there exist graph models yielding substantial improvements in both performance and efficiency compared to those generated by widely-used AR2N modeling [28].
@@ 3.2 Collection of Performance Metrics
For each task and each constructed graph model, we collected performance metrics, including training, validation, and test performance and runtime over at least five trials, and the parameter size of predictive models, which also varies with graph models. We provide the details below.
@@
Predictive Models. As predictive models, we used those provided by RelBench [28], which use Pytorch-Frame [16] to encode each table and pass the embeddings to graph neural networks (GNNs). Specifically, for classification and regression tasks, we used Heterogeneous GraphSAGE [13] with sum aggregation to update embeddings for the final predictions, and for recommendation tasks, we used ID-GNN [39]. As the RelBench models do not support edge modeling of tables, we extended them to incorporate edge features and used them when tables are modeled as edges (refer to Appendix A.2). We additionally used three alternative predictive models, as described in Section 4.5.
Training Details. Based on the training protocol (including the training, validation, and test splits) provided by RelBench [28]), we tuned only the learning rate for each graph model. The other hyperparameters were fixed to the combinations that yielded the best overall performance on 50 randomly sampled graph models. For classification and regression tasks, we selected the learning rate from {0.005, 0.001, 0.0005}, and for recommendation tasks, from {0.001, 0.0005, 0.0001}. The number of training epochs was fixed at 20, which was empirically sufficient for convergence. For reliability, we repeated each experiment with 15 different random seeds for the rel-f1 and rel-event datasets, and with 5 seeds for the remaining datasets. Refer to Appendix A.3 for further training details.
@@ 4 Observations from the RDB2G-Bench Datasets
Results. Figure 4 shows a representative case from the user-attendance task on the rel-event dataset. Note that all top-5 best performing graph models include the foreign-key relationship events $\rightarrow$ users. They all also model either the event_attendees table or the event_interest table as edges.
@@
Figure 3: Modeling table rows as edges (Row2Edge) can be crucial, depending on the task (Obs 2). It significantly improves performance for the user-repeat task, but not for the user-ignore task, even when both are defined on the same RDB.
Figure 4: Top-performing graph models share common substructures (Obs 3). As shown in their network schemas, the top-5 graph models commonly (a) include the foreign-key (FK) relationship events $\rightarrow$ users and (b) model either the event_attendees table or the event_interest table as edges. Note that the users_friends table has two FKs (user and friend) both referencing the users table.
@@
Implications. These results suggest that an effective RDB-to-graph modeling strategy should account not only for the characteristics of the RDB but also for those of the specific downstream task. Moreover, reusing graph model is effective only across tasks with closely aligned objectives.
@@
4.5 Obs 5. “Effectiveness of graph models generalizes across predictive ML models.”
Analysis Overview. We analyze performance correlations across different predictive ML models applied to the same graph models to examine whether a graph model effective with one ML model (e.g., Heterogeneous GraphSAGE [13]) is also effective with others. To this end, as additional ML models, we consider two graph neural networks (GNNs) (GraphSAGE with mean aggregation [13] and GIN [37]) and a graph transformer (GPS [26]) For their details, refer to Appendix A.2. As in the previous analysis, we employ the Spearman rank correlation [32].
Results. As shown in Table 25, the correlations are generally high (above 0.7) in most cases, indicating that the effectiveness of graph models tends to generalize across different ML models. Notably, the graph transformer also exhibits strong correlations with GNNs, despite their architectural differences.
Implications. This cross–ML model generalizability of graph models suggests that RDB-to-graph modeling strategies trained and shown to work well using our benchmark, RDB2G-Bench, can be effective across various predictive ML models, highlighting the broad utility of RDB2G-Bench.
Figure 6: Effectiveness of graph models generalizes across predictive ML models (Obs 5). Spearman correlations between different predictive models—GraphSAGE with sum and mean aggregation (denoted as sum and mean), GIN, and GPS—are generally high (above 0.8) across the three RDBs (rel-avito, rel-trial, and rel-event).
@@ 5 Benchmark Results on the RDB2G-Bench Datasets
In this section, we review the benchmark results of nine RDB-to-graph modeling methods on our RDB2G-Bench datasets.
@@ 5.1 Baselines for RDB-to-Graph Modeling
We benchmark nine RDB-to-graph modeling methods, which are categorized into (a) heuristic-based methods, (b) action-driven search algorithms, and (c) an LLM-based approach. All methods are designed to select a graph model among those in our RDB2G-Bench datasets (see Section 3.1 for details on our graph model space).
@@
  • S8. Reinforcement Learning (RL): It applies a controller based on recurrent neural networks [31, 14] and trains it via a policy gradient descent approach [36, 3]. The RL model learns to select actions of any type at each step of the RDB-to-graph modeling process to optimize downstream-task performance.
@@
LLM-Based Method. Inspired by AutoG [7], we developed an LLM-based baseline (S9) that directly generates action sequences for RDB-to-graph modeling. For fair comparisons with the above action-based search algorithms, it operates over the same predefined set of action types. The key distinction from AutoG lies in the simplified prompt design. Prompts irrelevant due to differences in action spaces are omitted. At each step, an LLM is provided with a set of candidate actions and selects a sequence of actions expected to be most effective, based on its reasoning capabilities. Claude Sonnet-3.5 is used as the backbone LLM, and details of the prompt design for action selection are provided in Appendix F.

### v2

@@ Abstract
[…]
@@ 1 Introduction
A relational database (RDB) is a collection of data organized into multiple tables connected by shared keys. RDBs enable systematic and efficient management of related information through query languages such as SQL, and have been widely adopted across diverse industries, including finance [18], healthcare [34], and e-commerce [30]. This widespread use has led to the emergence of diverse machine learning applications built upon RDBs.
@@
One important consideration in graph-based approaches is the variety of ways to model RDBs as graphs. For example, a single table row representing a user transaction on an item can be modeled either as (1) a node representing the transaction or (2) an edge linking the corresponding user and item. Additionally, one may choose to include only a subset of tables (i.e., nodes) and FK relationships (i.e., edges), excluding those considered less relevant. We call the process of representing RDBs as graphs RDB-to-graph modeling, and refer to the resulting graph representation as a graph model.
@@
Despite the importance of intelligent RDB-to-graph modeling, research in this area remains in its early stages, with one main reason being the difficulty of evaluation. First, there are numerous possible graph models (see Figure 2(a) for examples from real-world RDBs), making exhaustive comparisons extremely expensive. Moreover, evaluating even a single graph model is computationally costly, as it typically involves training graph neural networks (GNNs) on a large graph.
@@
[…]
@@ 2.1 Relational Deep Learning
Early approaches to relational database (RDB) learning relied on feature engineering [19] (e.g., table joins), transforming an RDB into a single table to leverage tabular machine learning [5, 21, 2, 13]. Recent works on relational deep learning (i.e., deep learning on RDBs) [8, 31, 11, 12, 9] pioneered modeling RDBs as graphs, which facilitate the capture of cross-table dependencies. Building on these graph models, advanced methods [6, 42] have achieved improved predictive performance on RDBs.
@@ 2.2 Benchmarks on Relational Deep Learning
Distinctive property of our benchmark: Existing benchmarks provide RDBs and predictive tasks to compare machine learning (ML) methods, which are typically based on the same graph presentation of the RDBs. Our benchmark, RDB2G-Bench, however, is designed to evaluate RDB-to-graph modeling strategies, which provide graph models based on which various modeling methods perform.
Specifically, by offering precomputed performance metrics for 50k graph models, RDB2G-Bench enables comparing various graph modeling strategies without repeated GNN training (see Section 5.2 for efficiency gains).
@@ 2.3 Graph-based Modeling of RDBs
Our benchmark, RDB2G-Bench, enables the evaluation of various RDB-to-graph modeling methods, including both heuristic and LLM-based approaches (see Section 5.1 for the ten methods included in the benchmark), ultimately fostering the development of more advanced modeling strategies.
@@ 3 Dataset Design for RDB2G-Bench
In this section, we introduce the datasets provided in RDB2G-Bench, our benchmark for the RDB-to-graph modeling problem. The problem is defined as follows:
Problem Definition.
Definition 1 (RDB-to-Graph Modeling).
Given: A relational database $\mathcal{R}$, and a graph neural network $\mathcal{M}$ for a downstream task $\mathcal{T}$. To Find: the graph model $\mathcal{G}^{*}$ of $\mathcal{R}$ that maximizes the performance $\mathcal{P}$ of $\mathcal{M}$ trained and evaluated on $\mathcal{T}$, i.e., $\mathcal{G}^{*}=\arg\max\nolimits_{\mathcal{G}\in\mathcal{F}(\mathcal{R},\mathcal{T})}\mathcal{P}(\mathcal{M}(\mathcal{G}),\mathcal{T}),$ where $\mathcal{F}(\mathcal{R},\mathcal{T})$ is the set of possible graph models of $\mathcal{R}$ on $\mathcal{T}$, defined in Section 3.1; and example metrics of $\mathcal{P}$ are provided in Section 3.2.
Each dataset comprises graph models (i.e., graph-structured representations) of a given relational database (RDB) paired with downstream task performance metrics of graph neural networks (GNNs) (denoted by $\mathcal{M}$) trained on those graph models. We first describe the considered design space of graph models (Section 3.1), followed by the procedure for obtaining GNN performance metrics on each graph model (Section 3.2)
@@ 3.1 Construction of Graph Models
As discussed in Section 1, there are various ways to model each RDB as a graph, and the performance of GNNs depends heavily on the chosen graph model. In RDB2G-Bench, we consider the design choices in the following two steps of RDB-to-graph modeling, both of which have a significant impact on the downstream-task performance of GNNs, as shown in Section 4:
@@
That is, for each RDB $\mathcal{R}$ (e.g., rel-avito), our dataset consists of graph models that result from different combinations of design choices in Step 1 and Step 2, which we denote by $\mathcal{F}(\mathcal{R},\mathcal{T})$ in Definition 1 (refer to Figure 1 for examples). These graph models are used to train graph neural networks (GNNs) for predictive tasks (e.g., user-clicks) defined on the RDB (as described in Section 3.2), and to be used for this purpose, the graph models must satisfy several constraints. First, a valid graph model must select the task table, which the considered predictive task is defined on. Second, all selected tables should be connected to the task table via a path whose length does not exceed the number of GNN layers. Violating this constraint would result in some nodes being unreachable during message passing, leading to degraded performance. Third, (the rows of) a table can be modeled as edges only if the table has exactly two FKs and its primary key (PK) is not referenced by any FKs in other tables. Note that, for a table with more than two FKs, hyperedge modeling would be required, which is beyond our scope, and edge modeling of a table whose PK is referenced by FK renders those FK relationships unrepresentable. Also note that these constraints may lead to different graph model spaces for downstream tasks defined on the same RDB. As summarized in Figure 2(a), we constructed about 50k graph models spanning the aforementioned RDBs and downstream tasks; and evaluated their downstream task performance metrics, as described below.
(a) Summary of the RDB2G-Bench datasets, which cover 50k graph models in total.
RDB Task Name Type # Tables # Graph Performance Statistics
Models Best AR2N [29] Worst
rel-avito user-clicks (UC) classification 8 944 67.93 64.66 60.89
user-visits (UV) classification 944 66.33 65.97 59.83
ad-ctr (AC) regression 1304 0.039 0.040 0.044
user-ad-visit (UAV) recommendation 909 3.682 3.661 0.159
rel-event user-repeat (UR) classification 5 214 82.29 77.65 63.96
user-ignore (UI) classification 214 82.82 82.22 74.29
user-attendance (UA) regression 214 0.237 0.244 0.266
rel-f1 driver-dnf (DD) classification 9 722 74.56 73.14 67.40
driver-top3 (DT) classification 722 81.88 78.11 75.37
driver-position (DP) regression 722 3.831 3.913 4.171
rel-stack post-post-related (PPL) recommendation 7 7979 12.04 10.82 0.006
rel-trial study-outcome (SO) classification 15 36863 70.91 68.09 62.85
(b) Visualization of the RDB2G-Bench datasets on three predictive tasks.
@@
Figure 2: (a) We summarize the RDBs, tasks, and their associated graph models. For each classification, regression, and recommendation task, we collect AUC-ROC (%), MAE, and MAP (%), respectively, on each graph model. For each task, we report the performances on the best graph model, the worst model, and that given by AR2N modeling [29]. (b) For three tasks (driver-top3, user-attendance, post-post-related), we visualize the distribution of performances on the downstream task (Y-axis) across all graph models, along with training time per epoch (X-axis) and the parameter size of the graph neural network (indicated by color). Note that there exist graph models yielding substantial improvements in both performance and efficiency compared to those generated by widely-used AR2N modeling [29].
@@ 3.2 Collection of Performance Metrics
For each task and each constructed graph model, we collected the following performance metrics (i.e., $\mathcal{P}$ in Definition 1):
  • Predictive performance: The training, validation, and test performances on the downstream task.
  • Runtime: The total elapsed time per epoch over at least five trials.
  • Parameter size: The total number of learnable parameters of predictive GNNs, which depends on the graph model.
Specifically, we collected the performance metrics under the following setups:
@@
Predictive GNNs. As predictive machine learning models, we used GNNs provided by RelBench [29], which leverage PyTorch-Frame [17] to encode each table as input to the GNNs. Specifically, for classification and regression tasks, we employed Heterogeneous GraphSAGE [14] with sum aggregation to update embeddings for the final predictions. For recommendation tasks, we utilized ID-GNN [41]. As the RelBench GNNs do not support edge modeling of tables, we extended them to incorporate edge features and used them when tables are modeled as edges (refer to Appendix A.2). We additionally used three alternative predictive GNNs, as described in Section 4.5.
Training Details. Based on the training protocol provided by RelBench [29] (including the training, validation, and test splits), we tuned only the learning rate for each graph model. The other hyperparameters were fixed to the combinations that yielded the best overall performance on 50 randomly sampled graph models. For classification and regression tasks, we selected the learning rate from {0.005, 0.001, 0.0005}, and for recommendation tasks, from {0.001, 0.0005, 0.0001}. The number of training epochs was fixed at 20, which was empirically sufficient for convergence. For reliability, we repeated each experiment with 15 different random seeds for the rel-f1 and rel-event datasets, and with 5 seeds for the remaining datasets. Refer to Appendix A.3 for further details.
@@ 4 Observations from the RDB2G-Bench Datasets
Results. Figure 4 shows a representative case from the user-attendance task on the rel-event dataset. Note that all top-5 best performing graph models include the foreign-key relationship events $\rightarrow$ users. They also model either the event_attendees table or the event_interest table as edges.
@@
Figure 3: Modeling table rows as edges (Row2Edge) can be crucial, depending on the task (Obs 2). EA (event_attendees), EI (event_interest), and UF (user_friends) indicate the tables whose rows can be modeled as edges. Note that Row2Edge modeling improves performance for the user-repeat task, but not for the user-ignore task, even when both are defined on the same RDB.
Figure 4: Top-performing graph models share common substructures (Obs 3). As shown in their graph models, the top-5 graph models commonly (a) include the foreign-key (FK) relationship events $\rightarrow$ users and (b) model either the event_attendees table or the event_interest table as edges. Note that the users_friends table has two FKs (user and friend) both referencing the users table.
@@
Implications. These results suggest that an effective RDB-to-graph modeling strategy should account not only for the characteristics of the RDB but also for those of the specific downstream task. Moreover, reusing a graph model is effective only across tasks with closely aligned objectives.
@@
4.5 Obs 5. “Effectiveness of graph models generalizes across predictive GNNs.”
Analysis Overview. We analyze performance correlations across different predictive GNNs applied to the same graph models to examine whether a graph model effective with one GNN (e.g., Heterogeneous GraphSAGE [14]) is also effective with others. To this end, we consider three additional GNNs (GraphSAGE with mean aggregation [14] and GIN [39]) and a graph transformer (GPS [27]) For their details, refer to Appendix A.2. As in the previous analysis, we employ the Spearman rank correlation [33].
Results. As shown in Figure 6, the correlations are generally high (above 0.7) in most cases, indicating that the effectiveness of graph models tends to generalize across different GNNs. Notably, the graph transformer also exhibits strong correlations with GNNs, despite their architectural differences.
Implications. This cross–GNN generalizability of graph models suggests that RDB-to-graph modeling strategies trained and shown to work well using our benchmark, RDB2G-Bench, can be effective across various predictive GNNs, highlighting the broad utility of RDB2G-Bench.
Figure 6: Effectiveness of graph models generalizes across predictive GNNs (Obs 5). Spearman correlations between different predictive GNNs—GraphSAGE with sum and mean aggregation (denoted as sum and mean), GIN, and GPS—are generally high (above 0.8) across the three RDBs (rel-avito, rel-trial, and rel-event).
@@ 5 Benchmark Results on the RDB2G-Bench Datasets
In this section, we review the benchmark results of ten RDB-to-graph modeling methods on our RDB2G-Bench datasets.
@@ 5.1 Baselines for RDB-to-Graph Modeling
We benchmark ten RDB-to-graph modeling methods, which are categorized into (a) heuristic-based methods, (b) action-driven search algorithms, and (c) an LLM-based approach. All methods are designed to select a graph model among those in our RDB2G-Bench datasets (see Section 3.1 for details on our graph model space).
@@
  • S8. Reinforcement Learning (RL): It applies a controller based on recurrent neural networks [32, 15] and trains it via a policy gradient descent approach [38, 3]. The RL-based predictor learns to select actions of any type at each step of the RDB-to-graph modeling process to optimize downstream-task performance.
@@
LLM-Based Baselines. Inspired by AutoG [7], we developed the following two LLM-based baselines that generate action sequences for RDB-to-graph modeling using LLMs.
  • S9. LLM baseline (LLM): It directly generates action sequences that specify how a given RDB is expressed as a graph based on the given prompts. At each step, an LLM is provided with a set of candidate actions and selects a sequence of actions expected to be most effective, based on its reasoning capabilities. The key distinction from AutoG lies in the simplified prompt design. Prompts irrelevant due to differences in action spaces are omitted.
  • S10. LLM-CoT baseline (LLM-CoT): It applies Chain-of-Thought [36] prompt design to encourage complex reasoning in the LLM baseline during the action selection step.
For fair comparisons with the above action-based search algorithms, both baselines operate over the same predefined set of action types. Claude Sonnet-3.5 is used as the backbone LLM, and details of the prompt designs are provided in Appendix F.